The time spent waiting in the line is approximately normally distributed. The mean waiting time is 5 minutes and the standard deviation of the waiting time is 1 minute. Find the probability that a person will wait for more than 3 minutes. Round your answer to four decimal places.
Answer #1
for normal distribution z score =(X-μ)/σx
here mean=       μ= 5
std deviation   =σ= 1.000

  probability that a person will wait for more than 3 minutes.:

probability =P(X>3)=P(Z>(3-5)/1)=P(Z>-2)=1-P(Z<-2)=1-0.0228=0.9772